40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 20 02:12:35 2014, Jenne, Update, LSC, ALS X noise investigation ALS_outofloop_19May2013.pdfXbeatSaturate.png
    Reply  Wed May 21 13:20:34 2014, manasa, Update, LSC, ALS X noise from angular motion of mirrors ALSX_OLYerrITM.png
       Reply  Thu May 22 00:30:40 2014, manasa, Update, LSC, ALS X noise from angular motion of mirrors ALSX_OLPerrITM2.pngALSX_OLYerrITM.pngALSX_OLPerrETM.pngALSX_OLYerrETM2.png
Message ID: 9972     Entry time: Tue May 20 02:12:35 2014     Reply to this: 9983
Author: Jenne 
Type: Update 
Category: LSC 
Subject: ALS X noise investigation 

[Rana, Jenne]

We have looked at a few things that do and don't affect the out of loop noise of the ALS X beat, and found that cavity alignment and beatnote RF frequency had the strongest effects.

Possible causes of noise:

1.  Air currents from A/C or flowbench.  No effect

        * When table lid is on, turning on and off the flow bench air did not qualitatively change the out of loop beatnote time series signal.

2.  Scattered light from other beams hitting green PDH PD.  No effect.

        * There are a few spots of green light that are hitting the case of the PDH photodiode, but when I put an iris in place to block those spots, there was no change in the beatnote spectra.  This makes sense to me since none of those spots were close to hitting the diode itself. 

        * Rana did notice that the beam was not well centered on the PD, so he steered the beam onto the center of the diode.  Also, the PD is now tilted a little bit so that the reflection from the diode doesn't go back into the beam path.  Neither of these things had an effect that we noticed in the beatnote noise.

3.  Oplev laser light getting to PDH PD.  Not tested.

       * We don't see any red light over by the PDH PD, so we did not try turning off the oplev's laser to see if that had an effect, but we suspect that it is not the cause of our noise.

4.  Clipping of main IR / green beam on Xend table.  Not tested.

      * We should still go have a look at this, but we no longer think that this is the main cause of the elevated noise.

5.  Scattered light all over Xend table.  Not tested.

     * We should still work on dumping extraneous beams on the table, but we do not think that this is the main cause of the elevated noise.

     * Rana took some photos so that we can see how truely bad the situation is.

6.  Amplitude modulation dip in NPRO.  Not tested.

    * It is probably still a good idea to check this, in case the dip in the amplitude modulation has changed over the year or two since it was last measured, but we also don't think that this is the main problem.

7.  Check PDH servo.  Not done.

     * I think this is still on Q's long-term todo list, but we should give the PDH servos a once-over.

8.  Arm cavity longitudinal motion.  No effect.

     * While the Xarm was locked with IR, we put a line at 1.7 Hz with 325 counts into the ITMX position.  To keep lock, the ETM had to move as well.  When we turned on this line (and increased its amplitude up to the final value of 325 cts), we did not see any qualitative change in the beatnote time series noise.

9.  Arm cavity alignment.  Significant DC effect.

    * When the alignment of one of the arm cavity mirrors is changed, the DC value of the beatnote signal changes. 

           * ITMX moved in yaw, we see a 7kHz/15urad DC shift in the BEATX_FINE_PHASE_OUT_HZ time series.

          * ETMX moved in yaw, we see an 8kHz/5.5urad DC shift in the time series.  We aren't sure why this is about a factor of 3 times larger effect (same shift for smaller misalignment) than the ITM.

    * We want to do a Yuta-style analysis to see what the angle to length coupling looks like, so that we can measure the angular motion of our cavity mirrors and put the expected noise into our ALS noise budget.  Perhaps this will help us understand the low frequency difference between our in-loop beatnote error signal and our in-loop PDH error signals (red vs. maroon on the ALS noise budget posted above Pianosa). 

    * I've asked Manasa to take some transfer functions in the morning, so that we can start to have an idea of what is going on with this.

10.  Beatnote RF frequency.  Significant broadband effect.

     * We have found that when the Xarm beatnote is at low RF frequencies, the noise is high, and when the beatnote is at high RF frequencies, the noise is low! 

     * Low RF freqs are below about 40 MHz, while high RF freqs are above about 90 MHz.  This has not been tested for the Yarm.  Also, these are for the case of "temp slider up, beatnote up".  I have not checked if the same is true for the other side of the PSL frequency, although I don't have reason to believe that it would be.

     * Maybe we are saturating some amplifiers?  We need to check this out.  One thought that Den mentioned was the harmonics, and that perhaps they are causing trouble in the electronics.

     * Den is going to think about implementing a frequency divider so that we can directly digitize the beatnote signal. 

    * Here are spectra for different cases:

          ALS_outofloop_19May2013.pdf

        * And here is a spectrogram showing us going back and forth between the high and low noise states:

          XbeatSaturate.png

                     *  A:  First noticing that noise is good when RF frequency is high.

                     * B:  Not locked on TEM00 mode, so extra noisy.  Disregard.

                     * C:  Bad noise time.  Xbeat was 21 MHz (dark purple on DTT spectrum above), Ybeat was 118 MHz (sea green on DTT spectrum above).

                    * D:  Good noise time. Xbeat was 89 MHz (light purple on DTT plot), Ybeat was still 118MHz (turquoise on DTT plot).

                     * E:  Bad noise time.  Xbeat was 37.5 MHz, Ybeat was still 118 MHz.

                     * F:  Good noise time.  Xbeat was 113 MHz, Ybeat was still 118 MHz.

ELOG V3.1.3-